A particle A of mass m moves on a smooth horizontal table and is connected by a light

inextensible string passing through a smooth hole O in the table to a particle B of the

same mas; m, which moves along the vertical through O. Initially B is at rest and a is the

distance from A to O, moving with velocity "sqrt{ga\/3}" perpendicular to O A.

What is the kinetic energy of the system?


1
Expert's answer
2021-12-07T20:11:09-0500

"E=\frac{E_1+E_2}2,"

"E_1=mga,"

"E_2=\frac{mv^2}2,"

"E=\frac{mga}2+m\sqrt{\frac{ga}{12}}."